Consider the following reaction: CO(g)+2H2(g)⇌CH3OH(g) A reaction mixture in a 5.17 −L flask at a certain temperature contains 27.0 g CO and 2.35 g H2. At equilibrium, the flask contains 8.67 g CH3OH.

Calculate the equilibrium constant (Kc) for the reaction at this temperature.

Kc=?
Answer #1

28-019/mot . И, 2.01glol CAOM 32.01 9/nol. Shpto Cafulate he molas f eac 2 73 23-017/=t= otefCO O.967mel 2- п, 2-0169moh 8-64ntequtbai O.0524 0-or24 So = 0 -1865-0.0524 0.13YI M 2ApGL2Y = 0.1202 C2255 - Tresorof 26-2